cos4thita/cos2phi+ sin 4 thita/sin2phi=1    given


to prove 


cos4phi/cos2thita+ sin 4 phi/sin2thita

Grade:10

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:Hello student, please find answer to your question
\frac{cos^{4}\theta }{cos^{2}\phi } + \frac{sin^{4}\theta }{sin^{2}\phi } = 1
\frac{cos^{4}\theta.sin^{2}\phi + sin^{4}\theta .cos^{2}\phi }{cos^{2}\phi.sin^{2}\phi } = 1
\frac{cos^{4}\theta(1-cos^{2}\phi) + sin^{4}\theta .cos^{2}\phi }{cos^{2}\phi.sin^{2}\phi } = 1
\frac{cos^{4}\theta-cos^{2}\phi.cos^{4}\theta + sin^{4}\theta .cos^{2}\phi }{cos^{2}\phi.sin^{2}\phi } = 1
\frac{cos^{4}\theta+cos^{2}\phi(sin^{4}\theta -cos^{4}\theta) }{cos^{2}\phi.sin^{2}\phi } = 1
\frac{cos^{4}\theta+cos^{2}\phi(sin^{2}\theta -cos^{2}\theta).(sin^{2}\theta +cos^{2}\theta ) }{cos^{2}\phi.sin^{2}\phi } = 1
\frac{cos^{4}\theta+cos^{2}\phi(sin^{2}\theta -cos^{2}\theta) }{cos^{2}\phi.sin^{2}\phi } = 1
You can proceed in this way.
Please rewrite what i need to prove.